[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 20, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 21 02:46:19 UTC 2016


HSU Shift Log for Wed Jul 20 2016 14:19:50 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues: 12Z TAFB Dvorak fixes imported on time. FSU Superensemble was  
late for both Darby and Estelle at 12Z - we didn't see it before the  
advisory went out for either system.

N-AWIPS/AWIPS2 issues: Microwave data has not come into NAWIPS since ~ 10z.

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160721/d2ef97b7/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list